OpenStudy (shiree19):

I have 2 questions today. Both are simplify the expression problems. #1 - (-11z)^0 A: z B: 1 C: -11 D: -1 I think it is B. Am I right? #2 - 14^-4 A: 1/14^4 B: -56 C: 1/14 D: 1/14^-4 I have no idea how to figure this out. On both problems can you please explain how you did it so I can do it in the future on my own?

OpenStudy (shiree19):

Please help @SolomonZelman @hippo4 @beccaboo333 and anyone else who knows how to do these problems.

OpenStudy (solomonzelman):

Anything to the zeroth power (besides zero,) equals 1.

OpenStudy (shiree19):

ok. so for the first I do -11z * 1 = -11z but that's not a choice.

OpenStudy (solomonzelman):

Did you read what I said?

OpenStudy (shiree19):

Yes. Anything to the zeroth power (besides zero,) equals 1. oh so besides 0. anything times 0 = 0 right?

OpenStudy (shiree19):

so if I do -11z * 0 = 0z = z or does it equal 1?

OpenStudy (solomonzelman):

\(\Large\color{black}{ \bf b^0=1 ~~~~~~~when }\) \(\Large\color{black}{ \bf b≠0 }\) Makes sense ?

OpenStudy (solomonzelman):

So what do you get for number 1 ?

OpenStudy (shiree19):

A = z.

OpenStudy (shiree19):

I am between A and B.

OpenStudy (solomonzelman):

You aren't reading what I am saying, OR you have improperly small background for this problem.

OpenStudy (shiree19):

I am reading what you are saying I just don't understand all of it.

OpenStudy (solomonzelman):

Anything to the zeroth power is equal to zero, besides zero. What doesn't make sense about it?

OpenStudy (shiree19):

if anything equals zero besides zero then what does zero equal?

OpenStudy (shiree19):

I'm sorry if this is hard to explain to me. It's very hard for me to understand this kind of math.

OpenStudy (solomonzelman):

Anything to the zeroth power is equal to \(\Large\color{black}{ \bf ONE }\) , besides zero. sorry for the errr So the answer is 1

OpenStudy (shiree19):

so it's B. 0 = 1 what happened to the -11z though? thanks for answering the 2nd problem.